An insurance company issues a policy on a horse trailer under the following conditions: The replacement cost ($6,000) will be paid for a total loss. If it is not a total loss, but the damage is more than $3,000, then $2,700 will be paid. Nothing will be paid for damage costing $3,000 or less and of course nothing is paid out if there is no damage. The company estimates the probability of the first three events as 0.08, 0.19, and 0.45 respectively. The amount the company should charge if it wishes to make a profit of $70 above the expected amount paid out in a year is:

$637
$1,275
$1,063
$850
Answer #1

Ans:

Expected amount to be paid=6000*0.08+2700*0.19+0*0.45=993

Amount the company should charge=993+70=1063

So,correct option is 1063